# Revisiting Semi-Supervised Learning in the Era of Foundation Models

### v1

@@ 6.2 Scalability of Ensembling
VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) (Cherti et al., 2023; Ridnik et al., 2021) and two ViT-L (CLIP, DINOv2).
PEFT Following the same setup, we start with LoRA and AdaptFormer, then expand our selection with four additional PEFT: ConvPass (Jie & Deng, 2022), BitFit (Zaken et al., 2022), VPT-Deep (Jia et al., 2022b), and Fact-TT (Jie & Deng, 2023). These methods span a broad range of PEFT approaches, including selective-based, adapter-based, and prompt-based techniques.
We evaluate CLIP on the DTD N3 dataset, training with LoRA, and examine ensemble performance across different numbers of pseudo-labels, ranging from 1 to 6. To gain deeper insights into the pattern, we examine two ensemble sources: (1) PEFT: Ensembling pseudo-labels generated by models trained with different PEFT methods while using the same pretrained VFM (CLIP) and (2) VFMs: Ensembling pseudo-labels generated by different VFMs, all trained using the same PEFT method (LoRA).
For each experiment, we report the average over 5 runs, each with a different randomly sampled set of pseudo-labels. As shown in Figure 6, we summarize the results as follows.
Both PEFT and VFMs Help. Whether we ensemble the pseudo-labels generated across different PEFT or VFMs, the performance is consistently improved. This indicates that both PEFT and VFMs contribute to the diversity of pseudo-labels and help to improve the ensemble robustness.
Diminishing Return. The performance improvement diminishes as we increase the number of ensemble sources. This suggests that the performance gain from ensemble is not linearly proportional to the number of sources. In order to achieve a balance between performance and computational cost, we recommend using a moderate number of ensemble sources, around 2 to 4 in this case.
6.3 Mean Label Consistently Outperforms
Furthermore, as also shown in Figure 6, we investigate three ensemble strategies: (1) Mean Label (our proposed method); (2) Mean Logits; and (3) Mean Probabilities.
Based on our observations from the previous section, we find that the Mean Label strategy is the most effective among the three. It consistently outperforms the other two strategies across different ensemble sources and ensemble sizes. This suggests that the Mean Label approach better captures the diversity of pseudo-labels and maximizes the information utilization from the ensemble sources.
6.4 Effectiveness of Hyperparameter Tuning
RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
@@ 6.4 Effectiveness of Hyperparameter Tuning
Figure 6: Scaling analysis for ensembling with more PEFT methods and VFMs. The starting point is pseudo-labels generated by the model itself fine-tuned with LoRA. The performance improves as more diverse pseudo-labels are incorporated, but the gain diminishes as the number of ensemble sources increases.
6.5 Impact of Labeled-Only Initialization on SSL
Per discussion in Section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT. To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch. These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in Section 6.
Table 5: Performance comparison of starting from a model trained on labeled data and then fine-tuning using SoftMatch. The results are reported on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). For consistency, we fix LoRA with dimension 4 for all the training.
DTD N6 68.4% 64.0% 63.5%
Diabetic Retinopathy N80 44.9% 44.5% 43.9%
KITTI N10 54.0% 27.9% 50.1%
@@ 7 Conclusion
As the machine learning community increasingly standardizes pre-trained foundation models as backbones across various applications, it becomes crucial to assess whether machine learning algorithms developed prior to the foundation model era remain effective. Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting. To address this gap, we conduct a comprehensive and systematic empirical study. Our findings reveal that VFMs can still benefit from SSL but require a specifically tailored approach. In particular, we propose a simple yet highly effective self-training algorithm, enhanced by pseudo-labels generated through the integration of multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques. Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. Specifically, compared to prior work in semi-supervised learning and vision foundation models, we do not see additional negative ethical and societal impacts that would be created by our work.
@@ Appendix A Dataset Details
  • DTD (Cimpoi et al., 2014): The Describable Texture Dataset (DTD) is designed for texture pattern recognition. It includes a diverse collection of images from 47 distinct texture categories, offering rich and varied examples to evaluate texture feature extraction and classification methods under natural conditions.
  • SUN397 (Xiao et al., 2010): SUN397 is a large-scale scene recognition benchmark that covers 397 categories of both indoor and outdoor scenes. It provides a comprehensive testbed for scene classification algorithms.
  • Resisc45 (Cheng et al., 2017): Resisc45 is a remote sensing image dataset specifically created for scene classification tasks. It comprises thousands of images across 45 categories, representing diverse landscapes such as urban, rural, and industrial areas. This dataset is instrumental in developing and benchmarking algorithms for geographic information systems and environmental monitoring.
  • Diabetic-Retinopathy (Dugas et al., 2015): This dataset focuses on high-resolution retinal images used for detecting diabetic retinopathy. It includes images that span various stages of the disease, providing a robust resource for training and evaluating deep learning models aimed at early diagnosis and automated medical analysis in ophthalmology.
  • Clevr-Count (Johnson et al., 2016): Derived from the CLEVR family of datasets, Clevr-Count is centered on visual reasoning, particularly object counting tasks. Using synthetically generated images, it challenges models to accurately count objects within complex scenes, thereby assessing their ability to understand spatial relationships and compositional structures in visual data.
  • KITTI-Dist (Geiger et al., 2012): KITTI-Dist is based on the renowned KITTI benchmark and is tailored for depth estimation in vehicular contexts. It provides high-quality stereo images along with ground-truth depth maps, supporting research in autonomous driving, 3D reconstruction, and stereoscopic vision, and serves as a key resource for evaluating depth estimation algorithms.
@@
DTD 752 1880
SUN397 17401 21750
Resisc45 5040 6300
Retinopathy 9207 42670
Clevr-C 14000 15000
KITTI 1354 711
@@ B.1 Computation Details
Our experiments were conducted on a workstation equipped with eight NVIDIA RTX 6000 Ada GPUs, two AMD EPYC 9554 64-Core Processors, and 800GB of RAM. Additionally, we utilized NVIDIA Tesla V100, NVIDIA Tesla A100, and NVIDIA RTX H100 GPUs for certain experiments. All experiments were implemented using PyTorch (Paszke et al., 2019).
@@ B.2 SSL and Labeled-Only Baseline Details
For all the experiments, we employ AdamW (Loshchilov & Hutter, 2019) optimizer with a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2017) with warm-up period with $2.5\%$ of total iterations. We use a batch size of $32+32$ for all experiments, where the first $32$ corresponds to labeled data and the second $32$ corresponds to unlabeled data. We list the hyperparameter search spaces for the SSL, Labeled-Only, and ST settings in Table 7, including the drop path rate (dpr) (Huang et al., 2016), training augmentation (train-aug), LoRA dimension (lora_dim), adapter bottleneck size (adapter_bottleneck), weight decay, momentum, learning rate (lr), and number of epochs. For other unmentioned SSL algorithm related hyperparameters, we use the default values provided in the original papers.
@@
As a completion of Section 6, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in Section B.2.
@@ C.1 Hyperparameter Tuning Procedure
Tuning without Labels The primary challenge in tuning hyperparameters within an unsupervised setting lies in effectively adjusting them without access to labels. Inspired by the approach for standardizing hyperparameter tuning in Unsupervised Domain Adaptation (UDA) (Ericsson et al., 2023), we propose tuning the hyperparameters using $\phi_{\mathcal{D}_{v}}$, $\ell_{\mathcal{D}_{v}}$ and $\pi_{\mathcal{D}_{v}}$. Specifically, we define function $v$ as a validator:
@@
To get the best performance model, we select the model with lowest value in $A$. We put the validators we use and more detailed examples in Section C.2.
@@ C.2 Validators
We deploys 7 semi-supervised learning validators, including
RankMe Score (Garrido et al., 2023) We adapt RankMe Score to compute the feature matrix rank of the pre-trained data over both source and target data domains.
Adjusted Mutual Information (AMI) (Vinh et al., 2009) We adapt AMI to compute the adjusted mutual information between predicted and cluster labels of the validation set.
Adjusted Rand Index (ARI) (Halkidi et al., 2002) ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions. Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:
@@
V-Measure (Rosenberg & Hirschberg, 2007) Beside AMI, we also adapt V-Measure to measure the harmonic mean between homogeneity and completeness over the clustering labels and prediction.
Fowlkes-Mallows Index (FMI) (Fowlkes & Mallows, 1983) Measures the similarity between two clusterings, FMI is defined by:
@@
Calinski-Harabasz Index (CHI) (Caliński & Harabasz, 1974) CHI is an internal clustering measurement. Given $n$ sets of data point $\{x\}_{1}^{n}$ grouped into $k$ clusters $\{C\}_{1}^{k}$ with centroids $\{c\}_{i}^{k}$ and $\bm{c}$ is the overall centroid, CHI is defined by:
@@
Batch Nuclear-norm Minimization (BNM) (Cui et al., 2020) An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.

### v2

@@ 6.2 Scalability of Ensembling
Figure 5: Scaling analysis for ensembling more PEFT methods and VFMs. Starting from LoRA fine-tuning, performance improves with diverse pseudo-labels, but gain diminishes with more ensemble sources.
VFMs Beyond ViT-B (CLIP, DINOv2), we incorporate four more VFMs, for a total of six, including four ViT-B (CLIP, DINOv2, OpenCLIP, ImangeNet21k) cherti2023reproducible ; ridnik2021imagenet21kpretrainingmasses and two ViT-L (CLIP, DINOv2).
PEFT Following the same setup, we start with LoRA and AdaptFormer, then expand our selection with four additional PEFT: ConvPass jie2022convolutionalbypassesbettervision , BitFit zaken2022bitfitsimpleparameterefficientfinetuning , VPT-Deep jia2022visualprompttuning , and Fact-TT jie2023factfactortuninglightweightadaptation . These methods span a broad range of PEFT approaches, including selective-based, adapter-based, and prompt-based techniques.
We train CLIP with LoRA on DTD N3, generate pseudo-labels, and ensemble groups of 1–6 labels. We compare (1) PEFT: ensemble of pseudo-labels from models using different PEFT methods on the same CLIP backbone, and (2) VFMs: ensemble of pseudo-labels from different VFMs all fine-tuned with LoRA. Results are averaged over five runs with distinct random pseudo-label sets (see Figure 5).
Mean Label Consistently Outperforms. As also shown in Figure 5, we investigate three ensemble strategies: (1) Mean Label (our proposed); (2) Mean Logits; and (3) Mean Probabilities. We find that the Mean Label strategy is the most effective among the three. It consistently outperforms the other two strategies across different ensemble sources and ensemble sizes. This suggests that the Mean Label approach better captures the diversity of pseudo-labels and maximizes the information utilization from the ensemble sources.
Both PEFT and VFMs Help. Whether we ensemble the pseudo-labels generated across different PEFT or VFMs, the performance is consistently improved, indicating that both PEFT and VFMs contribute to the diversity of pseudo-labels and improve the ensemble robustness.
Diminishing Return. The performance improvement diminishes as we increase the number of ensemble sources, suggesting the performance gain from ensemble is not linearly scalable. In order to achieve a balance between performance and computational cost, we recommend using around 2 to 4 in this case.
Impact of Pseudo-Label Quality. When applying V-PET, one might wonder: if the performances of different VFMs vary significantly, can their pseudo-labels still be ensembled to boost downstream results? The answer is yes, we demonstrate the result in section B.6.
Figure 6: Ranking frequency across SSL methods by the proposed benchmark. The number in $(i,j)$ indicates the frequency of method $i$ is ranked $j$-th across 12 settings. The number in bracket indicates the average rank.
@@ 7 Conclusion
As the machine learning community increasingly standardizes pre-trained foundation models as backbones across various applications, it becomes crucial to assess whether machine learning algorithms developed prior to the foundation model era remain effective. Specifically, given that vision foundation models (VFM) are pre-trained on millions of images, it remains unclear whether they would still benefit from abundant unlabeled data in conjunction with limited labeled data in a semi-supervised learning (SSL) setting. To this end, we conduct a comprehensive and systematic empirical study, revealing that VFMs can still benefit from SSL but require a specifically tailored approach. In particular, we propose a simple yet highly effective self-training algorithm that ensembles pseudo-labels from multiple VFMs and parameter-efficient fine-tuning (PEFT) techniques. Overall, our study offers actionable insights into SSL with VFMs and lays the groundwork for more scalable and practical semi-supervised learning in the era of foundation models.
@@ Appendix A Dataset Details
  • DTD [16]: The Describable Texture Dataset (DTD) is designed for texture pattern recognition. It includes a diverse collection of images from 47 distinct texture categories, offering rich and varied examples to evaluate texture feature extraction and classification methods under natural conditions.
  • SUN397 [64]: SUN397 is a large-scale scene recognition benchmark that covers 397 categories of both indoor and outdoor scenes. It provides a comprehensive testbed for scene classification algorithms.
  • Resisc45 [14]: Resisc45 is a remote sensing image dataset specifically created for scene classification tasks. It comprises thousands of images across 45 categories, representing diverse landscapes such as urban, rural, and industrial areas. This dataset is instrumental in developing and benchmarking algorithms for geographic information systems and environmental monitoring.
  • Diabetic-Retinopathy [22]: This dataset focuses on high-resolution retinal images used for detecting diabetic retinopathy. It includes images that span various stages of the disease, providing a robust resource for training and evaluating deep learning models aimed at early diagnosis and automated medical analysis in ophthalmology.
  • Clevr-Count [38]: Derived from the CLEVR family of datasets, Clevr-Count is centered on visual reasoning, particularly object counting tasks. Using synthetically generated images, it challenges models to accurately count objects within complex scenes, thereby assessing their ability to understand spatial relationships and compositional structures in visual data.
  • KITTI-Dist [27]: KITTI-Dist is based on the renowned KITTI benchmark and is tailored for depth estimation in vehicular contexts. It provides high-quality stereo images along with ground-truth depth maps, supporting research in autonomous driving, 3D reconstruction, and stereoscopic vision, and serves as a key resource for evaluating depth estimation algorithms.
@@
DTD 752 1,880
SUN397 17,401 21,750
Resisc45 5,040 6,300
Retinopathy 9,207 42670
Clevr-C 14,000 15,000
KITTI 1,354 711
@@ B.1 Computation Details
Our experiments were conducted on a workstation equipped with eight NVIDIA RTX 6000 Ada GPUs, two AMD EPYC 9554 64-Core Processors, and 800GB of RAM. Additionally, we utilized NVIDIA Tesla V100, NVIDIA Tesla A100, and NVIDIA RTX H100 GPUs for certain experiments. All experiments were implemented using PyTorch [51].
@@ B.2 SSL and Labeled-Only Baseline Details
For all the experiments, we employ AdamW [43] optimizer with a cosine annealing learning rate scheduler [42] with warm-up period with $2.5\%$ of total iterations. We use a batch size of $32+32$ for all experiments, where the first $32$ corresponds to labeled data and the second $32$ corresponds to unlabeled data. We list the hyperparameter search spaces for the SSL, Labeled-Only, and ST settings in Table 5, including the drop path rate (dpr) [33], training augmentation (train-aug), LoRA dimension (lora_dim), adapter bottleneck size (adapter_bottleneck), weight decay, momentum, learning rate (lr), and number of epochs. For other unmentioned SSL algorithm related hyperparameters, we use the default values provided in the original papers.
@@
As a completion of subsection 6.1, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in subsection B.2.
@@ B.3 PEFT Experiment Details
B.4 Linear Probing with Different Classification Heads
We provide results of frozen backbone with different classification heads on our benchmark in Table 8.
DTD SUN397 Resisc45 Retinopathy Clevr-C KITTI
N $6$ $6$ $2$ $80$ $20$ $10$
CLIP-LINEAR 61.80% 63.70% 69.30% 35.90% 33.10% 51.10%
CLIP-MLP-96 63.09% 60.29% 67.89% 36.91% 38.11% 50.77%
CLIP-MLP-768 62.29% 62.52% 69.62% 39.67% 39.51% 48.10%
CLIP LoRA 64.10% 66.79% 72.05% 42.16% 50.80% 64.28%
DINOv2-LINEAR 66.70% 65.30% 52.40% 41.20% 30.50% 51.30%
DINOv2-MLP-96 66.97% 62.41% 50.30% 44.33% 36.37% 50.91%
DINOv2-MLP-768 67.98% 61.16% 52.06% 35.66% 35.21% 50.77%
DINOv2 LoRA 68.40% 66.78% 63.08% 48.02% 61.72% 71.73%
Table 8: Linear probing results compared to LoRA under various scenarios, where LINEAR denote a single linear layer and MLP-x denotes an MLP with a single hidden layer of size x. The results demonstrate that LoRA is a more effective approach for fine-tuning vision foundation models.
B.5 Impact of Labeled-Only Initialization on SSL
Table 9: Performance of starting from a labeled-only model then fine-tuning with SoftMatch. Results on DTD (N6), Retinopathy (N80), and KITTI (N10). LoRA dim=4 is fixed.
DTD N6 68.4% 64.0% 63.5%
Retinopathy N80 44.9% 44.5% 43.9%
KITTI N10 54.0% 27.9% 50.1%
Per discussion in section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT. To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch. These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in subsection 6.1.
B.6 The Effectiveness of VFMs Ensemble
Table 10: We use pseudo-labels from different source to train a CLIP-B model. Demonstrating when performing ensembling training, even weaker pseudo-labels will boost performance.
Model Test Accuracy
CLIP 56.17%
ViT-B-IN21k 25.27%
CLIP + ViT-B-IN21k 58.62%
To assess the benefit of pseudo-label ensembling across VFMs, even when their individual accuracies differ significantly, we performed an ablation study on the CLIP-B model. We trained three variants on different pseudo-label sets: (1) labels generated by CLIP itself, (2) labels from ViT-B-IN21k, a comparatively weaker VFM, and (3) the ensemble of both sources. Results in Table 10 show that, despite ViT-B-IN21k’s lower standalone accuracy, combining its labels with CLIP’s improves overall performance.
B.7 The Effectiveness of Hyperparameter Tuning
RankMe AMI ARI V-Measure FMI CHI BNM Random Ours
@@ C.1 Hyperparameter Tuning Procedure
Tuning without Labels The primary challenge in tuning hyperparameters within an unsupervised setting lies in effectively adjusting them without access to labels. Inspired by the approach for standardizing hyperparameter tuning in Unsupervised Domain Adaptation (UDA) [23], we propose tuning the hyperparameters using $\phi_{\mathcal{D}_{v}}$, $\ell_{\mathcal{D}_{v}}$, and $\pi_{\mathcal{D}_{v}}$. Specifically, we define function $v$ as a validator:
@@
To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section C.2.
@@ C.2 Validators
We deploy 7 semi-supervised learning validators, including
RankMe Score [26] We adapt RankMe Score to compute the feature matrix rank of the pre-trained data over both source and target data domains.
Adjusted Mutual Information (AMI) [63] We adapt AMI to compute the adjusted mutual information between predicted and cluster labels of the validation set.
Adjusted Rand Index (ARI) [29] ARI is a widely used metric in cluster analysis by measuring the similarity between two clustering solutions. Given $RI$ is the Rank Index between the true and predicted clustering, ARI is defined by:
@@
V-Measure [55] Beside AMI, we also adapt V-Measure to measure the harmonic mean between homogeneity and completeness over the clustering labels and prediction.
Fowlkes-Mallows Index (FMI) [24] Measures the similarity between two clusterings. FMI is defined by:
@@
Calinski-Harabasz Index (CHI) [9] CHI is an internal clustering measurement. Given $n$ sets of data point $\{x\}_{1}^{n}$ grouped into $k$ clusters $\{C\}_{1}^{k}$ with centroids $\{c\}_{i}^{k}$ and $\bm{c}$ is the overall centroid, CHI is defined by:
@@
Batch Nuclear-norm Minimization (BNM) [18] An unsupervised domain adaptation method aims to generate predictions that favor both diversion and confidence; BNM maximizes the nuclear norm of the prediction matrix in a batch and repurposes it as a validation criterion.
@@ C.3 Hyperparameter Tuning Results
We present complete hyperparameter tuning results in Table 12.